a gas occupies a volume of 10l at stp, what is its final temperature if its volume is sl at 2atm?


1
Expert's answer
2020-07-27T05:36:11-0400

Answer.

Let the final temperature be T K

According to ideal gas equation,

(P1 x V1)/T1 = (P2 x V2)/T2

At STP, P1 = 1atm, V1 = 10l, T1 =273K

P2 = 2atm, V2 = S l, T2 = T

So, Putting these values,

T2 = (P2 x V2 x T1)/(P1 x V1) = (1 x 10 x273)/(2 x S) = (1365/S)K (Ans.)